# The Future of Artificial Intelligence and the Mathematical and Physical Sciences (AI+MPS)

### v2

@@ Abstract:
As we were finalizing this document, the White House released ``Winning the Race: America's AI Action Plan.'' [^1] https://www.whitehouse.gov/wp-content/uploads/2025/07/Americas-AI-Action-Plan.pdf This action plan echoes many of the findings of this document, in particular its recommendations to: Encourage Open-Source and Open-Weight AI (Sec. 2.3), Empower American Workers in the Age of AI (Sec. 2.7), Invest in AI-Enabled Science (Secs. 2.1, 2.5, 2.6, & 3), Build World-Class Scientific Datasets (Sec. 2.3), Advance the Science of AI (Sec. 2.2), Invest in AI Interpretability, Control, and Robustness Breakthroughs (Sec. 2.2), and Build an AI Evaluations Ecosystem (Sec. 2.3).
@@ 1 Introduction
Today’s artificial intelligence (AI) revolution is fueled by decades of basic research across multiple disciplines in the Mathematical and Physical Sciences (MPS), which have also served as important catalysts for AI developments by providing challenging problems, datasets, and novel application contexts that stimulate innovation. The recent 2024 Nobel Prizes, for example, recognized pioneering work in the development of foundational methods underpinning modern AI technologies (Physics) [^2] https://www.nobelprize.org/prizes/physics/2024/press-release/ and the applications of AI tools for computational protein design and structure prediction (Chemistry). [^3] https://www.nobelprize.org/prizes/chemistry/2024/press-release/ Indeed, the collection and careful curation of data (often taking decades) by MPS researchers have enabled breakthroughs with AI.
@@
Of course, the nature of a revolution is that it is disruptive and not always cohesive. The goal of this document is show how coordinated efforts can help the MPS domains best capitalize on, and contribute to, the future of AI. We outline a strategic vision for AI+MPS in Sec. 1.1, which also provides a roadmap for the rest of this white paper. (An additional roadmap appears in Sec. 4, which highlights key considerations for funding agencies, educational institutions, and individual AI+MPS researchers.) Then, in Sec. 1.2, we establish a common understanding/language across MPS for what we mean when we talk about AI, and summarize the virtuous cycle of AI+MPS so far, setting a baseline for how AI is advancing MPS and how MPS is advancing AI. Finally, in Sec. 1.3, we consider the potential for transformational outcomes that could result from dedicated AI+Science research.
@@ 1.1 Strategic Vision for AI+MPS
By addressing their unique domain challenges, scientists have continuously expanded the practical capabilities of AI, both for scientific applications and more broadly. In surveying the MPS community, we found many commonalities across disciplines in their approaches to AI, which point to opportunities for collaboration. By working together to developing techniques and exploring research opportunities, we can make substantial progress advancing AI+MPS as a research domain in its own right. To achieve this, it will be important to bolster the AI+MPS field through funding, infrastructure, collaboration, and education, while also identifying focused efforts within domains where AI will have the biggest impact.
@@ 1.2.1 AI in the Context of MPS
  • Supervised versus unsupervised learning: Supervised learning seeks to train regression or classification models to predict data labels. Unsupervised learning operates on unlabeled data to extract patterns, clusters, or low-dimensional structure, and is helpful for data discovery and outlier detection, which can be critical to account for ``unknown-unknowns'' and provide a (relatively) unbiased approach to data exploration. Semi-supervised learning blends these concepts to operate on partially labeled data sets. Self-supervised learning operates on unlabeled data, but where the training paradigm creates its own ``supervisory signal,'' often by predicting one part of the data from another (e.g., masked language modeling), and has been an important paradigm in training large language models (LLMs) where the data corpora are too large to admit human labeling of all training instances. In weakly-supervised learning, coarse-grained or inaccurate labels are available, while in self-supervised learning, labels are automatically generated during pretest task training and transfer learning is applied to the downstream training task of interest.
@@
    Generative models aim to learn a process that can generate new data points that resemble a given set of samples, even though the true underlying distribution that produced the original data is unknown. Examples of generative models are generative adversarial networks (GANs), variational autoencoders (VAEs), autoregressive models, normalizing flows (NFs), and diffusion models. Recently, generative models have driven much of the progress across multiple domains and in multiple applications, including anomaly detection, surrogate modeling, and simulation-based inference (SBI) (Sec. 2.5.1).
@@
  • Large language models (LLMs): The term ``generative models'' has taken on prominence in the context of LLMs. Starting from a ``prompt'' as input, these algorithms output rich text responses that can describe scientific concepts, generate hypotheses, or even suggest experimental designs. While LLMs have demonstrated impressive capabilities, the debate continues about their fundamental nature. Some researchers still maintain they can be viewed as ``stochastic parrots'' that primarily recombine training data in sophisticated ways. Others see them as a potential pathway toward ``artificial general intelligence'' (AGI) , pointing to their ability to reason about novel scientific questions and integrate knowledge across domains. Genuine successes in this area have been reliant upon logic-based reasoning engines (e.g., Lean [^4] https://lean-lang.org for formal math proofs) to perform reasoning with the rigor typically required in MPS, while LLMs have been limited to translating into appropriate syntax and/or invoking such logic engines as agents.
@@
  • Symbolic AI and formal methods: Logic-based symbolic methods have traditionally played a crucial role in computer science, particularly in specifying and verifying hardware, software, and communication protocols. These formal methods primarily utilize Boolean satisfiability problems (SATs) and more generalized frameworks such as satisfiability modulo theories (SMT). Fundamental computational breakthroughs in these areas during the 1990s significantly improved the scalability and practicality of symbolic reasoning tools, according to a report by the National Academies. [^5] https://nap.nationalacademies.org/catalog/27241/artificial-intelligence-to-assist-mathematical-reasoning-proceedings-of-a-workshop In contemporary mathematics, automated reasoning has become a central focus, leading to the development of first-order theorem provers and interactive proof assistants, enabling machine-assisted proofs for problems such as the Pythagorean triplet problem. Recently, integrating symbolic methods with AI has spurred innovations in mathematical reasoning systems, such as AlphaProof and AlphaGeometry, which demonstrate unprecedented capabilities in automated theorem proving and geometric reasoning.
@@ 1.2.2 How AI is Accelerating Scientific Discovery in MPS
  • Accelerating simulations: Scientific machine learning (SciML) techniques accelerate simulations in various scientific and engineering domains by addressing the limitations of traditional computational methods, which is discussed further in Sec. 2.5.2. AI can be used to predict the time evolution of complex dynamical systems and construct efficient surrogate models that can make predictions at a fraction of the computational cost, which is especially critical in outer-loop and many-query problems in inference and design. The learned collective variables can also expose new fundamental understandings of the dynamical systems and the models can be profitably constructed within virtuous iterative loops of simulation and model training. Algorithms leveraging the function approximation properties of neural networks have also enabled simulation of high-dimensional dynamical systems beyond reach of traditional techniques.
@@ 1.2.3 How MPS Research is Driving AI Understanding
  • Quantifying uncertainty and ensuring reliability: In an effort to ensure AI-generated scientific predictions are robust, reliable, and suitable for high-stakes scientific applications, MPS researchers are developing rigorous techniques for UQ, error estimation, and out-of-distribution detection that have the potential to be extremely important for many other application domains, particularly those where it is much more difficult to recognize unreliable predictions. Sec. 2.5.3 discusses UQ opportunites in more detail.
@@
  • Providing a safe ecosystem of development: Many MPS fields offer an ideal sandbox for developing AI tools in an environment where the data and problems are less likely to lead directly to negative human impacts. Such a safe environment could be the key to developing AGI, without being severely red-taped by policy. Sec. 3 provides insight into the domain-specific problems that can be used for developing AI and their impact on AI innovation. While safe, most of the MPS discoveries are high stakes, which are important for national prestige and competitiveness on the world stage.
@@ 1.3 The Future of AI+MPS
There are many opportunities across domains to advance AI+MPS and in doing so, have a substantial impact on research outcomes in both AI and MPS. Though it is challenging to predict what would be the most impactful, we can extrapolate some possibilities based on existing efforts. Current AI advances are trending towards the development of more capable AI co-scientists and could even lead to fully autonomous AI scientists that can generate complex, scientifically meaningful hypotheses, design experiments, and interpret results. Ideally, such AI co-scientists would operate in a way that preserves (and even enhances) existing standards for scientific rigor and integrity. Perhaps the most ambitious goal is the Nobel Turing challenge, [^6] https://doi.org/10.1038/s41540-021-00189-3 wherein AI makes a discovery at the Nobel level. These are certainly interesting and potentially worthwhile pursuits, but they would require substantial domain-specific meta-analysis on the use of such tools.
@@ 2.1 Advocate for Diverse Funding Streams
  • Short-term to long-term projects: Short-term projects may be sufficient for some advances, especially if they have clearly established goals or are led by individuals with interdisciplinary expertise, but long-term projects may be more effective in cases where researchers need to build collaboration and/or are exploring opportunities to advance the field. Small awards with rapid review times enable nimble research, especially due to the exceedingly fast pace of development in AI. For example, programs like NSF ACCESS [^7] https://access-ci.org have enabled quick turnaround times for small requests for computing resources (though funding for computing is typically separate from funding for people). At the other end of the spectrum, longer-term funding on the 5–10 year scale can better enable the development of large-scale AI projects and ensure their sustainability for the community.
For any of these AI+MPS funding streams, proposal evaluation will need to take into account both cross-cutting and domain-specific issues. For example, it would help support innovative research to ensure that the evaluation criteria for reviewers not only consider the expertise of the principal investigatorss (PIs) in each of the involved disciplines, but also their ability to assemble and lead an interdisciplinary team, which involves a complementary set of skills that are not always considered in evaluation metrics. For interdisciplinary proposals, it may be (and is often) the case that the variety of expertise required to evaluate the full scope of the proposed research cannot be covered by individual reviewers, and thus multiple individuals may be required to evaluate the work in full. This creates additional overhead to consolidate input, reach a consensus, and decide how to weigh strengths/weaknesses of AI aspects against MPS aspects. Assessment of individual interdisciplinary grants may benefit from having ad hoc review panels to supplement traditional panels that collectively review a full set of submissions.
@@ 2.1.1 Institute-Scale Activities
To support interdisciplinary work in AI, researchers will need funding to support community hubs and research collaboration. Currently, the NSF and DOE fund high-profile centers, such as the NSF National AI Research Institutes, [^8] https://aiinstitutes.org Centers for Chemical Innovation [^9] https://www.nsf.gov/mps/che/centers-chemical-innovation such as the Center for Computer Assisted Synthesis that attracted substantial industrial funding, and big collaborations led by the national labs, which have shown to be successful in building cross-disciplinary teams, leading to impactful results. There are also examples of collaborations with funding partners; for example, the Simons Foundation has been equal partners with the NSF on both Center- and Institute-level AI-focused/related projects in MPS, including the Mathematical and Scientific Foundations of Deep Learning (MoDL) collaboratives (2020), the National Institute for Theory and Mathematics in Biology (NITMB) (2023), and most recently the two new AST-themed National AI Research Institutes (2024). Broadening beyond this base is essential to grow a larger and more vibrant AI+Science community. One way to approach this is to leverage the existing institutes and build on the work they have already done, such as sharing educational and community resources, developing best practices for AI+Science, and supporting infrastructure. Combining efforts rather than launching new efforts every few years would allow the community to build on the momentum of existing institutes.
@@ 2.1.2 Project-Scale Activities
An example of a program at the NSF that supports project-based funding for software development and scientific domain research is the Cyberinfrastructure for Sustained Scientific Innovation (CSSI) program. [^10] https://www.nsf.gov/funding/opportunities/cssi-cyberinfrastructure-sustained-scientific-innovation This could serve as a model for AI+MPS funding, with the software aspect of CSSI replaced by AI. The model co-funds grants between the Directorate for Computer and Information Science and Engineering (CISE) and a science domain, e.g. MPS/PHY. The solicitation includes software-specific questions and information, e.g. deployment plans, long-term maintenance, etc., but is open to all participating domains. The reviewers in this program are specifically chosen to be experts at software development, and come from various domains. Projects are only funded if the science domain is willing to co-fund, which ensures that the domain application is exciting, and the reviewers being experts on the technical side ensures that the right people are chosen to be funded to perform the project. Of course, in such a case, it would be important for the expectations that will be reviewed by both domain researchers and AI experts to be clearly stated in the initial solicitations.
@@ 2.1.3 Individual Investigators
Currently, researchers have few avenues to obtain funding that allows them to focus their time on bridging gaps between MPS domain science and AI. One promising approach is to support a dedicated connector role—a ``centaur'' or polymath—by offering incentives for universities to train or hire individuals in this capacity. For example, establishing a three-year Investigator program for PIs who have a proven track record in interdisciplinary initiatives could empower them to effectively link diverse teams and mentor emerging researchers. An existing example of this is the Schmidt Science Polymaths Program, [^11] https://www.schmidtsciences.org/schmidt-science-polymaths/ which funds recently-tenured professors working on interdisciplinary research broadly, including but not restricted to AI+Science. This could also lead to opportunities for expanding retraining and up-skilling programs (see more in Sec. 2.7.1), which would enable established scientists from both the domain and AI sides to immerse themselves in the complementary discipline, ensuring that more senior researchers remain agile and well-equipped to drive innovative, cross-disciplinary collaborations. Ideally, this type of funding would include a foreseeable pathway for individual researchers from smaller institutions (like R2 and smaller) to develop more comprehensive and theoretical AI research. For example, funding faculty to buy them out of their teaching for 6–12 months so that they can pursue opportunities for upskilling would complement the current 3-year Investigator program, helping to scale to a broader community of researchers.
A similar approach could be used for early-career researchers. Funding postdoctoral and/or graduate student fellowships at the intersection of AI and domain science would attract interdisciplinary experts from the next generation of talent who have already begun to gain experience connecting communities (see Secs. 2.7.2 & 2.7.3). There have been examples of this as a successful model in programs like the NSF's MPS-ASCEND program, [^12] https://www.nsf.gov/funding/opportunities/mps-ascend-mathematical-physical-sciences-ascending-postdoctoral the NSF IAIFI Fellowship, [^13] https://iaifi.org/fellows.html and the Schmidt AI in Science Postdoctoral Fellowship. [^14] https://schmidtsciencefellows.org Importantly, researchers hired into these roles should have substantial independence in their research direction, allowing them to pursue innovative ideas and collaborate effectively across domains.

### v3

@@ Abstract:
As we were finalizing this document, the White House released ``Winning the Race: America's AI Action Plan'' [undef]. This action plan echoes many of the findings of this document, in particular its recommendations to: Encourage Open-Source and Open-Weight AI (Sec. 2.3), Empower American Workers in the Age of AI (Sec. 2.7), Invest in AI-Enabled Science (Secs. 2.1, 2.5, 2.6, & 3), Build World-Class Scientific Datasets (Sec. 2.3), Advance the Science of AI (Sec. 2.2), Invest in AI Interpretability, Control, and Robustness Breakthroughs (Sec. 2.2), and Build an AI Evaluations Ecosystem (Sec. 2.3).
@@ 1 Introduction
Today’s artificial intelligence (AI) revolution is fueled by decades of basic research across multiple disciplines in the Mathematical and Physical Sciences (MPS), which have also served as important catalysts for AI developments by providing challenging problems, datasets, and novel application contexts that stimulate innovation. The recent 2024 Nobel Prizes, for example, recognized pioneering work in the development of foundational methods underpinning modern AI technologies (Physics) [undefa] and the applications of AI tools for computational protein design and structure prediction (Chemistry) [undefb]. Indeed, the collection and careful curation of data (often taking decades) by MPS researchers have enabled breakthroughs with AI.
@@
Of course, the nature of a revolution is that it is disruptive and not always cohesive. The goal of this document is to show how coordinated efforts can help the MPS domains best capitalize on, and contribute to, the future of AI. We outline a strategic vision for AI+MPS in Sec. 1.1, which also provides a roadmap for the rest of this white paper. (An additional roadmap appears in Sec. 4, which highlights key considerations for funding agencies, educational institutions, and individual AI+MPS researchers.) Then, in Sec. 1.2, we establish a common understanding/language across MPS for what we mean when we talk about AI, and summarize the virtuous cycle of AI+MPS so far, setting a baseline for how AI is advancing MPS and how MPS is advancing AI. Finally, in Sec. 1.3, we consider the potential for transformational outcomes that could result from dedicated AI+Science research.
@@ 1.1 Strategic Vision for AI+MPS
By addressing their unique domain challenges, scientists have continuously expanded the practical capabilities of AI, both for scientific applications and more broadly. In surveying the MPS community, we found many commonalities across disciplines in their approaches to AI, which point to opportunities for collaboration. By working together to develop techniques and explore research opportunities, we can make substantial progress toward advancing AI+MPS as a research domain in its own right. To achieve this, it will be important to bolster the AI+MPS field through funding, infrastructure, collaboration, and education, while also identifying focused efforts within domains where AI will have the biggest impact.
@@ 1.2.1 AI in the Context of MPS
  • Supervised versus unsupervised learning: Supervised learning seeks to train regression or classification models to predict data labels. Unsupervised learning operates on unlabeled data to extract patterns, clusters, or low-dimensional structure, and is helpful for data discovery and outlier detection, which can be critical to account for ``unknown-unknowns'' and provide a (relatively) unbiased approach to data exploration. Semi-supervised learning blends these concepts to operate on partially labeled data sets. Self-supervised learning operates on unlabeled data, but where the training paradigm creates its own ``supervisory signal,'' often by predicting one part of the data from another (e.g., masked language modeling). It has been an important paradigm in training large language models (LLMs) where the data corpora are too large to admit human labeling of all training instances. In weakly-supervised learning, coarse-grained or inaccurate labels are available, while in self-supervised learning, labels are automatically generated during pretest task training and transfer learning is applied to the downstream training task of interest.
@@
    Generative models aim to learn a process that can generate new data points that resemble a given set of samples, even though the true underlying distribution that produced the original data is unknown. Examples of generative models are generative adversarial networks (GANs), variational autoencoders (VAEs), autoregressive models, normalizing flows (NFs), and diffusion models [undefc]. Recently, generative models have driven much of the progress across multiple domains and in multiple applications, including anomaly detection, surrogate modeling, and simulation-based inference (SBI) (Sec. 2.5.1).
@@
  • Large language models (LLMs): The term ``generative models'' has taken on prominence in the context of LLMs. Starting from a ``prompt'' as input, these algorithms output rich text responses that can describe scientific concepts, generate hypotheses, or even suggest experimental designs. While LLMs have demonstrated impressive capabilities, the debate continues about their fundamental nature. Some researchers still maintain they can be viewed as ``stochastic parrots'' that primarily recombine training data in sophisticated ways. Others see them as a potential pathway toward ``artificial general intelligence'' (AGI) , pointing to their ability to reason about novel scientific questions and integrate knowledge across domains. Genuine successes in this area have been reliant upon logic-based reasoning engines (e.g., Lean for formal math proofs [undefd]) to perform reasoning with the rigor typically required in MPS, while LLMs have been limited to translating into appropriate syntax and/or invoking such logic engines as agents.
@@
  • Symbolic AI and formal methods: Logic-based symbolic methods have traditionally played a crucial role in computer science, particularly in specifying and verifying hardware, software, and communication protocols. These formal methods primarily utilize Boolean satisfiability problems (SATs) and more generalized frameworks such as satisfiability modulo theories (SMT). Fundamental computational breakthroughs in these areas during the 1990s significantly improved the scalability and practicality of symbolic reasoning tools, according to a report by the National Academies [undefe]. In contemporary mathematics, automated reasoning has become a central focus, leading to the development of first-order theorem provers and interactive proof assistants, enabling machine-assisted proofs for problems such as the Pythagorean triplet problem. Recently, integrating symbolic methods with AI has spurred innovations in mathematical reasoning systems, such as AlphaProof and AlphaGeometry, which demonstrate unprecedented capabilities in automated theorem proving and geometric reasoning.
@@ 1.2.2 How AI is Accelerating Scientific Discovery in MPS
  • Accelerating simulations: Scientific machine learning (SciML) techniques [undeff] accelerate simulations in various scientific and engineering domains by addressing the limitations of traditional computational methods, as discussed further in Sec. 2.5.2. AI can be used to predict the time evolution of complex dynamical systems and construct efficient surrogate models that can make predictions at a fraction of the computational cost, which is especially critical in outer-loop and many-query problems in inference and design. The learned collective variables can also expose new fundamental understandings of the dynamical systems and the models can be profitably constructed within virtuous iterative loops of simulation and model training. Algorithms leveraging the function approximation properties of neural networks have also enabled simulation of high-dimensional dynamical systems beyond reach of traditional techniques.
@@ 1.2.3 How MPS Research is Driving AI Understanding
  • Quantifying uncertainty and ensuring reliability: In an effort to ensure AI-generated scientific predictions are robust, reliable, and suitable for high-stakes scientific applications, MPS researchers are developing rigorous techniques for UQ, error estimation, and out-of-distribution detection that have the potential to be extremely important for many other application domains, particularly those where it is much more difficult to recognize unreliable predictions [undefg]. Sec. 2.5.3 discusses UQ opportunites in more detail.
@@
  • Providing a safe ecosystem of development: Many MPS fields offer an ideal sandbox for developing AI tools in an environment where the data and problems are less likely to lead directly to negative human impacts. Such a safe environment could be the key to developing AGI, without being severely red-taped by policy. Sec. 3 provides insight into the domain-specific problems that can be used for developing AI and their impact on AI innovation. While safe, most of the MPS discoveries are high stakes and important for national prestige and competitiveness on the world stage.
@@ 1.3 The Future of AI+MPS
There are many opportunities across domains to advance AI+MPS and in doing so, have a substantial impact on research outcomes in both AI and MPS. Though it is challenging to predict what would be the most impactful, we can extrapolate some possibilities based on existing efforts. Current AI advances are trending toward the development of more capable AI co-scientists and could even lead to fully autonomous AI scientists that can generate complex, scientifically meaningful hypotheses, design experiments, and interpret results. Ideally, such AI co-scientists would operate in a way that preserves (and even enhances) existing standards for scientific rigor and integrity. Perhaps the most ambitious goal is the Nobel Turing challenge [undefh], wherein AI makes a discovery at the Nobel level. These are certainly interesting and potentially worthwhile pursuits, but they would require substantial domain-specific meta-analysis on the use of such tools.
@@ 2.1 Advocate for Diverse Funding Streams
  • Short-term to long-term projects: Short-term projects may be sufficient for some advances, especially if they have clearly established goals or are led by individuals with interdisciplinary expertise, but long-term projects may be more effective in cases where researchers need to build collaboration and/or are exploring opportunities to advance the field. Small awards with rapid review times enable nimble research, especially due to the exceedingly fast pace of development in AI. For example, programs like NSF ACCESS [undefi] have enabled quick turnaround times for small requests for computing resources (though funding for computing is typically separate from funding for people). At the other end of the spectrum, longer-term funding on the 5–10 year scale can better enable the development of large-scale AI projects and ensure their sustainability for the community.
For any of these AI+MPS funding streams, proposal evaluation will need to take into account both cross-cutting and domain-specific issues. For example, it would help support innovative research to ensure that the evaluation criteria for reviewers not only consider the expertise of the principal investigators (PIs) in each of the involved disciplines, but also their ability to assemble and lead an interdisciplinary team, which involves a complementary set of skills that are not always considered in evaluation metrics. For interdisciplinary proposals, it may be (and is often) the case that the variety of expertise required to evaluate the full scope of the proposed research cannot be covered by individual reviewers, and thus multiple individuals may be required to evaluate the work in full. This creates additional overhead to consolidate input, reach a consensus, and decide how to weigh strengths/weaknesses of AI aspects against MPS aspects. Assessment of individual interdisciplinary grants may benefit from having ad hoc review panels to supplement traditional panels that collectively review a full set of submissions.
@@ 2.1.1 Institute-Scale Activities
To support interdisciplinary work in AI, researchers will need funding to support community hubs and research collaboration. Currently, the NSF and DOE fund high-profile centers, such as the NSF National AI Research Institutes [undefj], Centers for Chemical Innovation [undefk] like the Center for Computer Assisted Synthesis that attracted substantial industrial funding, and big collaborations led by the national labs, which have shown to be successful in building cross-disciplinary teams, leading to impactful results. There are also examples of collaborations with funding partners; for example, the Simons Foundation has been equal partners with the NSF on both Center- and Institute-level AI-focused/related projects in MPS, including the Mathematical and Scientific Foundations of Deep Learning (MoDL) collaboratives (2020), the National Institute for Theory and Mathematics in Biology (NITMB) (2023), and most recently the two new AST-themed National AI Research Institutes (2024). Broadening beyond this base is essential to grow a larger and more vibrant AI+Science community. One way to approach this is to leverage the existing institutes and build on the work they have already done, such as sharing educational and community resources, developing best practices for AI+Science, and supporting infrastructure. Combining efforts rather than launching new efforts every few years would allow the community to build on the momentum of existing institutes.
@@ 2.1.2 Project-Scale Activities
An example of a program at the NSF that supports project-based funding for software development and scientific domain research is the Cyberinfrastructure for Sustained Scientific Innovation (CSSI) program [undefl].This could serve as a model for AI+MPS funding, with the software aspect of CSSI replaced by AI. The model co-funds grants between the Directorate for Computer and Information Science and Engineering (CISE) and a science domain, e.g. MPS/PHY. The solicitation includes software-specific questions and information, e.g. deployment plans, long-term maintenance, etc., but is open to all participating domains. The reviewers in this program are specifically chosen to be experts at software development, and come from various domains. Projects are only funded if the science domain is willing to co-fund, which ensures that the domain application is exciting, and the reviewers being experts on the technical side ensures that the right people are chosen to be funded to perform the project. Of course, in such a case, it would be important for the expectations that will be reviewed by both domain researchers and AI experts to be clearly stated in the initial solicitations.
@@ 2.1.3 Individual Investigators
Currently, researchers have few avenues to obtain funding that allows them to focus their time on bridging gaps between MPS domain science and AI. One promising approach is to support a dedicated connector role—a ``centaur'' or polymath—by offering incentives for universities to train or hire individuals in this capacity. For example, establishing a three-year Investigator program for PIs who have a proven track record in interdisciplinary initiatives could empower them to effectively link diverse teams and mentor emerging researchers. An existing example of this is the Schmidt Science Polymaths Program [undefm], which funds recently-tenured professors working on interdisciplinary research broadly, including but not restricted to AI+Science. This could also lead to opportunities for expanding retraining and up-skilling programs (see more in Sec. 2.7.1), which would enable established scientists from both the domain and AI sides to immerse themselves in the complementary discipline, ensuring that more senior researchers remain agile and well-equipped to drive innovative, cross-disciplinary collaborations. Ideally, this type of funding would include a foreseeable pathway for individual researchers from smaller institutions (like R2 and smaller) to develop more comprehensive and theoretical AI research. For example, funding faculty to buy them out of their teaching for 6–12 months so that they can pursue opportunities for upskilling would complement the current 3-year Investigator program, helping to scale to a broader community of researchers.
A similar approach could be used for early-career researchers. Funding postdoctoral and/or graduate student fellowships at the intersection of AI and domain science would attract interdisciplinary experts from the next generation of talent who have already begun to gain experience connecting communities (see Secs. 2.7.2 & 2.7.3). There have been examples of this as a successful model in programs like the NSF's MPS-ASCEND program [undefn], the NSF IAIFI Fellowship [undefo], and the Schmidt AI in Science Fellowship [undefp]. Importantly, researchers hired into these roles should have substantial independence in their research direction, allowing them to pursue innovative ideas and collaborate effectively across domains.